# Unraveling the Single Tangent Space Fallacy: An Analysis and Clarification for Applying Riemannian Geometry in Robot Learning

### v2

@@ II-A Riemannian Manifolds
A Riemannian manifold $(\mathcal{M},g)$ is a smooth manifold endowed with a Riemannian metric $g$, that is a family of inner products $g_{\bm{x}}:\mathcal{T}_{\bm{x}}\mathcal{M}\times\mathcal{T}_{\bm{x}}\mathcal{M}\rightarrow\mathbb{R}$ associated to each point $\bm{x}\in\mathcal{M}$ [9]. To operate with Riemannian manifolds, it is common practice to exploit their Euclidean tangent spaces and resort to mappings back and forth between $\mathcal{T}_{\bm{x}}\mathcal{M}$ and $\mathcal{M}$. Specifically, the exponential map $\text{Exp}_{\bm{x}}\left(\bm{u}\right):\mathcal{T}_{\bm{x}}\mathcal{M}\to\mathcal{M}$ maps a point $\bm{u}\in\mathcal{T}_{\bm{x}}\mathcal{M}$ to a point $\bm{y}$ on the manifold, so that it lies on the geodesic starting at $\bm{x}$ in the direction $\bm{u}$ and such that the geodesic distance $d_{\mathcal{M}}(\bm{x},\bm{y})=d_{\mathbb{R}}(\bm{x},\bm{u})$ (see Fig. 2-right). The inverse operation is the logarithmic map $\text{Log}_{\bm{x}}\left(\bm{y}\right):\mathcal{M}\to\mathcal{T}_{\bm{x}}\mathcal{M}$ Finally, the parallel transport $\Gamma_{\bm{x}\rightarrow\bm{y}}(\bm{u}):\mathcal{T}_{\bm{x}}\mathcal{M}\to\mathcal{T}_{\bm{y}}\mathcal{M}$ describes how elements of $\mathcal{M}$ can be transported along curves on $\mathcal{M}$ while maintaining their intrinsic geometric properties. This allows us to operate elements lying on different tangent spaces (see Fig. 2-right).
@@ III-B Euclidean GMM on a Single Tangent Space
Figure 3: Datapoints and GMM of Figs. 0(c)0(d) in the single tangent spaces at the origin with a top view (left) and Fréchet mean (right).
@@ IV-A A Deep Dive into Mathematics of Tangent Spaces
$a[\bm{\gamma}_{1}]+b[\bm{\gamma}_{2}]=d\varphi_{\bm{x}}^{-1}\left(a\varphi_{\bm{x}}([\bm{\gamma}_{1}])+b\varphi_{\bm{x}}([\bm{\gamma}_{2}])\right).$ (7)
@@ IV-B Fallacies of Single Tangent Space Approaches
The exponential is a local diffeomorphism.

### v3

@@ II-A Riemannian Manifolds
A Riemannian manifold $(\mathcal{M},g)$ is a smooth manifold endowed with a Riemannian metric $g$, that is a family of inner products $g_{\bm{x}}:\mathcal{T}_{\bm{x}}\mathcal{M}\times\mathcal{T}_{\bm{x}}\mathcal{M}\rightarrow\mathbb{R}$ associated to each point $\bm{x}\in\mathcal{M}$ [9]. To operate with Riemannian manifolds, it is common practice to exploit their Euclidean tangent spaces and resort to mappings back and forth between $\mathcal{T}_{\bm{x}}\mathcal{M}$ and $\mathcal{M}$. Specifically, the exponential map $\text{Exp}_{\bm{x}}\left(\bm{u}\right):\mathcal{T}_{\bm{x}}\mathcal{M}\to\mathcal{M}$ maps a point $\bm{u}\in\mathcal{T}_{\bm{x}}\mathcal{M}$ to a point $\bm{y}$ on the manifold, so that it lies on the geodesic starting at $\bm{x}$ in the direction $\bm{u}$ and such that the geodesic distance $d_{\mathcal{M}}(\bm{x},\bm{y})=d_{\mathbb{R},g_{\bm{x}}}(\bm{x},\bm{u})$ (see Fig. 2-right). The inverse operation is the logarithmic map $\text{Log}_{\bm{x}}\left(\bm{y}\right):\mathcal{M}\to\mathcal{T}_{\bm{x}}\mathcal{M}$ Finally, the parallel transport $\Gamma_{\bm{x}\rightarrow\bm{y}}(\bm{u}):\mathcal{T}_{\bm{x}}\mathcal{M}\to\mathcal{T}_{\bm{y}}\mathcal{M}$ describes how elements of $\mathcal{M}$ can be transported along curves on $\mathcal{M}$ while maintaining their intrinsic geometric properties. This allows us to operate elements lying on different tangent spaces (see Fig. 2-right).
@@ III-B Euclidean GMM on a Single Tangent Space
Figure 3: Datapoints and GMM of Figs. 1(c)1(d) in the single tangent spaces at the origin with a top view (left) and Fréchet mean (right).
@@ IV-A A Deep Dive into Mathematics of Tangent Spaces
$a[\bm{\gamma}_{1}]+b[\bm{\gamma}_{2}]=d\varphi_{\bm{x}}^{-1}\left(a\;\!d\varphi_{\bm{x}}([\bm{\gamma}_{1}])+b\;\!d\varphi_{\bm{x}}([\bm{\gamma}_{2}])\right).$ (7)
@@ IV-B Fallacies of Single Tangent Space Approaches
The exponential map is a local diffeomorphism.